OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)).

OpenStudy (anonymous):

g(-1) = the value of g when x = -1, so g(-1) = (-1)^2 -7(-1) - 9 = 1+7-9 = -1 Now, f(g(-1)) is just f(-1) = -1-2 = -3 Final answer is -3.

OpenStudy (anonymous):

g(-1)=(-1)2-7(-1)-9=4 f(-4)=-4-2=-6 final answer is -6